OpenStudy (anonymous):

FancyUsername you are my favourite math expert. Just saying XD 7x+4y=39 2x+4y=14 A. (5, 4) B. (3, 4) C. (3, 2) D. (5, 1) I really hope the next exam is a bit different and easy so I may know how to do it and I won't be asking so many questions. XD

OpenStudy (anonymous):

you use the elimination method. and substract those equations. 5x = 25 x = 25/5 = 5 2 * 5 + 4y = 14 4y = 14 - 10 = 4 y = 1 check: 7 * 5 + 4 = 39 35 + 4 = 39 39 = 39 Variant D(5 1)